OpenStudy (murkwundg):

Which polynomial has factors of 4x – 7 and x + 4? 3x2 + x – 3 4x2 + 9x – 28 3x2 – 7x – 3 4x2 – 23x – 28

OpenStudy (queenie):

(4x - 7)(x + 4)....solve by using FOIL

OpenStudy (queenie):

do u know the FOIL method ?

OpenStudy (murkwundg):

no

OpenStudy (queenie):

(4x - 7)(x + 4) FOIL F -- multiply first terms in each set....4x * x = O -- multiply outside terms...4x * 4 = ? I -- multiply inside terms...-7 * x = ? L -- multiply last terms..-7 * -4 = ? now do ur multiplying and what do u get ?

OpenStudy (queenie):

so whats 4x * x

OpenStudy (queenie):

its 4x^2 and whats 4x * 4.....its 16x and -7 * x = -7x and -7 * 4 = -28 so we have : 4x^2 + 16x - 7x - 28....combine like terms 4x^2 + 9x - 28 <-- ur answer

OpenStudy (murkwundg):

appreciate it